Let z = cosθ + isinθ.
Then z^n = cos(nθ) + isin(nθ) for all n ∈ N(by de Moivre) and z^−n = cos(nθ)−isin(nθ).

(a) Show that 2cos(nθ) = z^n + z^−n and 2isin(nθ) = z^n −z^−n

(b) Show that 2^ncos^n θ = (z + 1/z)^n and (2i)^n sin^n θ = (z− 1 z/n)

(c) Use (b) to express sin^7 θ in terms of multiple angles

(d) Express cos^3 θsin^4 θ in terms of multiple angles

(e) Eliminate θ from the equations 4x = cos(3θ) + 3cosθ; 4y = 3sinθ−sin(3θ).
1
Expert's answer
2020-05-13T18:41:08-0400

"\\bold{Solution} \\\\\n\n\\bold {a)}\\\\\n\nz^{n} + z^{-n} = cos(n\u03b8)+ i sin(n\u03b8) + cos(n\u03b8) - i sin(n\u03b8) = 2cos(n\u03b8) \\\\\n\nz^{n} - z^{-n} = cos(n\u03b8)+ i sin(n\u03b8) - cos(n\u03b8) + i sin(n\u03b8) =2isin(n\u03b8) \n\\\\\n\\bold{b)} \\\\\n\nFrom \\ a)\\ follow \\ that \\ when \\ n=1 \\\\\n\nz^{1} + z^{\u22121}= 2cos(\u03b8)\\\\\n\n( z^1 + z^{\u22121})^n =(2cos(\u03b8) )^n = 2^n cos^n(\u03b8)\\\\ \n\nz^1 - z^{\u22121}= 2i sin(\u03b8)\\\\\n\n( z^1 - z^{\u22121})^n =(2i sin(\u03b8) )^n = (2i)^n sin^n(\u03b8) \\\\\n\n\n\\bold{c) }\\\\\n\nFrom \\ b) \\ we\\ have \\ that\\\\\n\n(2i sin\u03b8)^7 = (z-z^{-1})^7 =\nC_{0 }^7 z^7 (-(z^{-1}))^0 + C_{1}^7 z^6 (-(z^{-1}))^1 + C_{2}^7 z^5 (-(z^{-1}))^2+C_{3}^7 z^4 (-(z^{-1}))^3+ C_{4}^7 z^3 (-(z^{-1}))^4 + C_{5}^7 z^2 (-(z^{-1}))^5 +C_{6}^7 z^1 (-(z^{-1}))^6 +C_{7}^7 z^0 (-(z^{-1}))^7 =\\\\\n\n= 1z^7 - 7z^6 z^{-1} + 21z^5 z^{-2} - 35z^4 z^{-3} +35z^3 z^{-4} - 21z^2 z^{-5} + 7z^1 z^{-6}-1z^0 z^{-7} =\\\\\n\n= z^7 \u2013 7 z^5+ 21 z^3 \u2013 35z + 35 z^{-1} \u2013 21z^{-3} + 7z^{-5} - z^{-7}=\\\\\n\n= ( z^7 - z^{-7}) \u2013 7( z^5 - z^{-5}) +21( z^3-z^{-3} ) \u2013 35(z - z^{-1} ) =\\\\\n\n=2isin\u20617\u03b8 \u2013 7*2isin\u20615\u03b8 + 21* 2isin\u20613\u03b8 \u2013 35*2isin\u2061\u03b8 =\\\\\n\n= 2i (sin\u20617\u03b8 \u2013 7sin\u20615\u03b8 + 21sin\u20613\u03b8 \u2013 35sin\u2061\u03b8)\\\\\n\n\n\n(2i)^7 (sin\u03b8)^7 =(2i sin\u03b8)^7\\\\\n\nsin^7 \u03b8=\\frac{sin\u20617\u03b8 \u2013 7 sin\u20615\u03b8 + 21 sin\u20613\u03b8 \u2013 35 sin\u2061\u03b8}{-64}\\\\\n\n\n\n\\bold{d)}\\\\\n\n2^3 cos^3 \u03b8 = (z^1+ z^{-1})^3 = 1z^3 + 3z^2 z^{-1} + 3 z^1 z^{-2} + 1z^0 z^{-3} =\\\\\n\n= z^3+3z+3z^{-1}+ z^{-3} =\\\\\n\n= (z^3+ z^{-3} ) +3(z^1+ z^{-1}) =\\\\\n\n= 2cos\u20613\u03b8 + 3 *2cos\u2061\u03b8 =\\\\\n\n= 2cos\u20613\u03b8 + 6cos\u2061\u03b8\\\\\n\ncos^3 \u03b8 = \\frac{1}{4} cos\u20613\u03b8 + \\frac{3}{4} cos\u2061\u03b8\\\\\n\n(2i)^3 sin^4 \u03b8 = (z^1-z^{-1})^4 = 1z^4 + 4z^3 (-z^{-1})^1 + 6 z^2 (-z^{-1})^2 + 4 z^1 (-z^{-1})^3 + 1 (-z^{-1})^4 =\\\\\n\n= z^4-4z^2 +6 \u2013 4 z^{-2} + z^{-4} =\\\\\n\n= (z^4+ z^{-4})-4(z^2+ z^{-2})+ 6 =\\\\\n\n= 2cos\u20614\u03b8 \u2013 4*2cos2\u2061\u03b8+6\\\\\n\nsin^4 \u03b8= \\frac{1}{8} cos\u20614\u03b8- \\frac{1}{2} cos2\u2061\u03b8+\\frac{3}{8}\\\\\n\n\ncos^3 \u03b8 sin^4 \u03b8= (\\frac{1}{4} cos\u2061(3\u03b8) + \\frac{3}{4} cos\u2061(\u03b8))(\\frac{1}{8} cos\u2061(4\u03b8)- \\frac{1}{2} cos(2\u2061\u03b8)+\\frac{3}{8}) =\\frac{cos(-\\theta)+cos(7\\theta)}{64} - \\frac{cos(\\theta)+cos(5\\theta)}{16} + \\frac{3}{32}cos(3\\theta) + \\frac{3(cos(-3\\theta)+cos(5\\theta))}{64} - \\frac{3(cos(-\\theta)+cos(3\\theta))}{16} +\\frac{9}{32}cos(\\theta)=3cos(\\theta)-3cos(3\\theta)-cos(5\\theta)+cos(7\\theta)\\\\\n\n\n\n\n\n\\bold{e)}\\\\ \n\nFrom \\ a)\\ we\\ have\\ that\\\\\n\n4x = cos\u20613\u03b8 + 3 cos\u2061\u03b8\\\\\n\n4y = 3sin\u2061\u03b8 - sin\u20613\u03b8\\\\\n\ncos\u20613\u03b8 = \\frac{z^3+ z^{-3}}{2} = \\frac{(z^1+ z^{-1})(z^2-z z^{-1}+ z^{-2}))}{2} = \\\\\n=\\frac{2 cos\u2061\u03b8 (2 cos2\u2061\u03b8-1))}{2} = cos\u2061\u03b8(4cos^2 \u03b8 \u2013 3) =\\\\\n= 4cos^3 \u03b8 \u2013 3cos\u2061\u03b8\\\\\n\n4x = 4cos^3 \u03b8\\\\\n\nx = cos^3 \u03b8\\\\\n\ncos\u2061\u03b8=\\sqrt[3]{x}\\\\\n\\theta=arccos(\\sqrt[3]{x})\\\\\n\n\n\nsin\u20613\u03b8 = \\frac{z^3- z^{-3}}{2} = \\frac{(z^1- z^{-1})(z^2+z z^{-1}+ z^{-2})}{2i} = \\\\\n=\\frac{2i sin\u2061\u03b8 (2 cos2\u2061\u03b8+1)}{2i}= sin\u2061\u03b8(3 - 4sin^2 \u03b8 ) = 3sin\u2061\u03b8 - 4sin^3 \u03b8\\\\\n\n4y = 4sin^3 \u03b8\\\\\n\ny = sin^3 \u03b8\\\\\n\nsin\u2061\u03b8=\\sqrt[3]{y}\\\\\n\\theta=arcsin(\\sqrt[3]{y})\\\\"